# Africa-Centric Self-Supervised Pre-Training for Multilingual Speech Representation in a Sub-Saharan Context

### v2

@@ Abstract
We present the first self-supervised multilingual speech model trained exclusively on African speech. The model learned from nearly 60 000 hours of unlabeled speech segments in 21 languages and dialects spoken in sub-Saharan Africa. On the SSA subset of the FLEURS-102 dataset, our approach based on a HuBERT (0.09B) architecture shows competitive results, for ASR downstream task, compared to the w2v-bert-51 (0.6B) pre-trained model proposed in the FLEURS benchmark, while being more efficient by using 7x less data and 6x less parameters. Furthermore, in the context of a LID downstream task, our approach outperforms FLEURS baselines accuracy by over 22%.
@@ 2 Datasets
Conneau et al. (2022) publicly released a parallel speech dataset in 102 languages and proposed it as benchmark. Data are divided in seven macro family, including a sub-Saharan Africa group. We therefore evaluate our approach on this SSA subset (FLEURS) which is composed of 20 languages, 5 of which are present in our pre-trained dataset.
@@ 3 Experiments
For downstream task training, we used the SpeechBrain toolkit (Ravanelli et al., 2021). The final pretrained model is considered as a speech encoder and is fully fine-tuned with two 1024 linear layers and a softmax output at the top. A first pool of speech recognition system (60k) is obtained by a direct fine-tuning of the whole model on each language of the FLEURS dataset. A second pool (60k) is then obtained by first jointly fine-tuning on all languages before fine-tuning again on each language.
Following the methodology of the FLEURS paper (Conneau et al., 2022) and to be consistent with their results, we did not rescore the hypothesis with a language model. Average character error rates (CERs) obtained on the 20 languages of the FLEURS test set are given in table 1. The detailed scores per language are provided in appendix B.
@@
To ensure the quality of the speech representation, we fine-tuned our pretrained model using SpeechBrain for a language identification (LID) downstream task. We employed adaptive average pooling to produce output with shape [Batch,1,20] and we applied a softmax. We call this model 60K. The model is trained for 15 epochs on the 20 languages of the FLEURS subset.
We also propose a second scenario where we employed adaptive average pooling to produce output with shape [Batch,1,768], with the addition of two linear layers to smoothly decrease the dimension from 768 to 256 then from 256 to 20. We call this model 60K. It is trained under the same conditions as 60K. Accuracy for both scenarios is presented in Table 2.
| FLEURS | FLEURS | 60K | 60K | |
| FLEURS | 59.1 | 62.2 | 84.9 | 90.4 |

### v3

@@ Abstract
We present the first self-supervised multilingual speech model trained exclusively on African speech. The model learned from nearly 60 000 hours of unlabeled speech segments in 21 languages and dialects spoken in sub-Saharan Africa. On the SSA subset of the FLEURS-102 dataset, our approach based on a HuBERTbase (0.09B) architecture shows competitive results, for ASR downstream task, compared to the w2v-bert-51 (0.6B) pre-trained model proposed in the FLEURS benchmark, while being more efficient by using 7x less data and 6x less parameters. Furthermore, in the context of a LID downstream task, our approach outperforms FLEURS baselines accuracy by over 22%.
@@ 2 Datasets
Conneau et al. (2022) publicly released a parallel speech dataset in 102 languages and proposed it as benchmark. Data are divided in seven macro family, including a sub-Saharan Africa group. We therefore evaluate our approach on this SSA subset (FLEURSSSA) which is composed of 20 languages, 5 of which are present in our pre-trained dataset.
@@ 3 Experiments
For downstream task training, we used the SpeechBrain toolkit (Ravanelli et al., 2021). The final pretrained model is considered as a speech encoder and is fully fine-tuned with two 1024 linear layers and a softmax output at the top. A first pool of speech recognition system (60k(0.09B)) is obtained by a direct fine-tuning of the whole model on each language of the FLEURS dataset. A second pool (60kFT-ALL(0.09B)) is then obtained by first jointly fine-tuning on all languages before fine-tuning again on each language.
Following the methodology of the FLEURS paper (Conneau et al., 2022) and to be consistent with their results, we did not rescore the hypothesis with a language model. Average character error rates (CERs) obtained on the 20 languages of the FLEURSSSA test set are given in table 1. The detailed scores per language are provided in appendix B.
@@
To ensure the quality of the speech representation, we fine-tuned our pretrained model using SpeechBrain for a language identification (LID) downstream task. We employed adaptive average pooling to produce output with shape [Batch,1,20] and we applied a softmax. We call this model 60KLID. The model is trained for 15 epochs on the 20 languages of the FLEURSSSA subset.
We also propose a second scenario where we employed adaptive average pooling to produce output with shape [Batch,1,768], with the addition of two linear layers to smoothly decrease the dimension from 768 to 256 then from 256 to 20. We call this model 60KLID-smooth. It is trained under the same conditions as 60KLID. Accuracy for both scenarios is presented in Table 2.
FLEURSw2v-bert | FLEURSmSLAM | 60KLID | 60KLID-smooth | |
FLEURSSSA | 59.1 | 62.2 | 84.9 | 90.4 |